A car with a mass of 2500 kg and moving at 3 m/s northeast collides with a 3000 kg car moving northwest at 2.5 m/s . Find their final velocities. Assume a perfectly inelastic collision.


1
Expert's answer
2021-03-23T08:15:05-0400

Let's choose the axes as shown in the figure below.



In a perfectly inelastic collision the momentum conservation law still holds. Thus, the total momentum before the collision in the projection on the x-axis is:


"p_{x} = -m_1v_1"

is equal to the total momentum after the collision in the projection on the x-axis:


"p_{x}' = (m_1 + m_2)v_x'"

where "m_1 = 2500kg, m_2 = 3000kg" are the masses of the cars, "v_1 = 3m\/s, v_2 = 2.5m\/s" are their speeds before the collision, and "v_x'" is their speed after the collsion in x-direction. Thus, obtain:


"-m_1v_1 = (m_1 + m_2)v_x'\\
v_x' = -\dfrac{m_1v_1}{m_1 + m_2} = -\dfrac{2500\cdot 3}{2500+3000} \approx -1.36m/s"

Similarly, the momentum conservation law in the projection on y-axis gives:


"m_2v_2 = (m_1 + m_2)v_y'"

where "v_y'" is car's speed after the collsion in y-direction. Thus, obtain:


"m_2v_2 = (m_1 + m_2)v_y'\\
v_y' = -\dfrac{m_2v_2}{m_1 + m_2} = -\dfrac{3000\cdot 2.5}{2500+3000} \approx 1.36m/s"

Thus, the final velocity is:


"\mathbf{v}' = (-1.36, 1.36)\space m/s"

Answer. "\mathbf{v}' = (-1.36, 1.36)\space m/s".